OpenStudy (anonymous):

write an equvilant logarithmic equation from each exponential equation 3 exponent 3 = 27.................5 negative exponent 2 = 1/25

OpenStudy (anonymous):

\[3^3 = 27\]\[\log_{3}27 = 3 \]\[5^{-2} = 1/25\]\[\log_{5}(1/25) = -2 \]

OpenStudy (anonymous):

ok sooo thanks but can you explain please

OpenStudy (anonymous):

so i do it in exponential its going in reverse

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Here's the rule (you need to memorize this: if: \[ a^b = c\] then:\[\log_{a}c = b \]

OpenStudy (anonymous):

ok thaks and if i do it exponential its puting back to were the equation is to be solved as in logarithmic